OpenStudy (anonymous):

Given the linear equation y = -3x – 2, find the y-coordinates of the points (0, ), (-3, ), (-1, ). Please show all of your work. Plot these points and graph the linear equation.

OpenStudy (anonymous):

You're given the x-values that you need (they're the first numbers in the ordered pairs you have). So, for example, the first x-value is 0, so you plug this into the equation to find the y-value: y = -3(0)-2 = -2 So (0,-2) is the first point. You do this for the other two, then plot them. Do you know how to plot points?

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

y=-3(-3)-2 = 9 - 2 = 7 Second point: (-3,7) y = -3(-1)-2 = 3-2 = 1 Third point (-1,1).

OpenStudy (anonymous):

You draw a set of 'axes' like the one I'm going to send in an attachment.

OpenStudy (anonymous):

The horizontal line is the 'x-axis' and the vertical line is the 'y-axis'. To plot each of the points, you look at the first number in the ordered pair and count along the x-axis from the origin the number of units indicated. Next look at the second number in the ordered pair and move the number of units indicated, but this time, move along the y-axis. So for the point, (-3,7), you'll see I've counted -3 from the origin, and then moved up +7. Where you stop is the point you plot.

OpenStudy (anonymous):

To graph the equation, you just draw a line through the three points.

OpenStudy (anonymous):

Is this okay?

OpenStudy (anonymous):

i got it now thanks

OpenStudy (anonymous):

np :)